# DSPy Assertions: Computational Constraints for Self-Refining Language Model Pipelines

### 20 Dec 2023 v1

@@ 2.2 Motivating Example
3 Semantics of LM Assertions
To help with the goals mentioned above, in this work, we introduce LM Assertions and integrate them in DSPy. We define LM Assertions as programmatic elements that dictate certain conditions or rules that must be adhered to during the execution of an LM pipeline. These constraints ensure that the pipeline’s behavior aligns with specified invariants or guidelines, enhancing the reliability, predictability, and correctness of the pipeline’s output.
Delineating Assert from Conventional Assertions.
The conventional assert statement, built into most programming languages, is a debugging aid that checks a condition and, if the condition evaluates to false, raises an AssertionError exception, typically terminating the program execution. In contrast, our Assert construct offers a sophisticated retry mechanism, while supporting a number of other new optimizations. On an Assert failing, the pipeline transitions to a special retry state, allowing it to reattempt a failing LM call while being aware of its previous attempts and the error message raised. If, after a maximum number of self-refinement attempts, the assertion still fails, the pipeline transitions to an error state and raises an AssertionError, terminating the pipeline. This enables Assert to be much more powerful than conventional assert statements, leveraging the LM to conduct retries and adjustments before concluding that an error is irrecoverable.
Delineating Suggest from Assert.
In contrast to our Assert statements, our Suggest statements are softer constraints that recommend but do not mandate conditions that may guide the LM pipeline toward desired domain-specific outcomes. When a Suggest condition is not met, similar to Assert, the pipeline enters the special retry state, allowing reattempts of the failing LM call and self-refinement. However, if the suggestion fails after a maximum number of self-refinement attempts, the pipeline simply logs a warning SuggestionError message and continues execution. This allows the pipeline to adjust its behavior in response to the suggestion while being flexible and resilient to suboptimal states (or sub-optimal or heuristic computational checks).
In the following sections, we define the default backtracking semantics of these constructs more formally. However, we provide the opportunity to extend custom semantics for both Assert and Suggest (Section 4.2).
3.1 Semantics of Assert
The Assert construct enforces invariants within the LM pipeline. The semantics of an assertion can be defined in terms of a state transition system where $\sigma_{r}$ represents the pipeline’s state, and the subscript $r$ represents the current retry count within the state $\sigma$. The maximum number of retries allowed per assertion is denoted by $R$. The transitions can be defined as follows:
$\sigma_{r}\vdash\mathtt{Assert}(e,m)$ $\rightarrow\sigma_{r+1}\quad$
if $\text{if }\text{eval}(\sigma,e)=\text{false}\text{ and }r<R$
$\sigma_{r}\vdash\mathtt{Assert}(e,m)$ $\rightarrow\sigma^{\bot}\quad$
if $\text{if }\text{eval}(\sigma,e)=\text{false}\text{ and }r\geq R$
3.2 Semantics of Suggest
$\sigma_{r}\vdash\mathtt{Suggest}(e,m)$ $\rightarrow\sigma_{r+1}\quad$
if $\text{if }\text{eval}(\sigma,e)=\text{false}\text{ and }r<R$
$\sigma_{r}\vdash\mathtt{Suggest}(e,m)$ $\rightarrow\sigma_{0}^{\prime\prime}\quad$
if $\text{if }\text{eval}(\sigma,e)=\text{false}\text{ and }r\geq R$
3.3 Handling Self-Refinement
From the above, both Assert and Suggest allow the pipeline to retry a failing LM call and self-refine its outputs in a special retry state. One might observe that this involves dynamically altering the control flow of the LM pipeline during execution. On passing assertions and suggestions, the control flows typically into the next LM pipeline module. However, to handle failures, the LM pipeline may define an error handler that determines the next instruction to execute. The handler takes the current erring state $\sigma$ and the error message $m$ as inputs and returns a new state. In the new state, control flows as described in Section 3.1 and 3.2. For both Assert and Suggest, if the maximum retry attempts are not surpassed, the handler yields the control to the failing LM module with an updated prompt that includes past failing outputs and instructions. However, upon exceeding the maximum retries, the handler halts the execution for a failing Assert or progresses to the next module in the pipeline for a Suggest.
In the next section, we describe the implementation of these constructs and handlers in DSPy.
4 Implementation
4.1 API Design
dspy.Assert(constraint: bool, msg: Optional[str],
                backtrack: Optional[module])
dspy.Suggest(constraint: bool, msg: Optional[str],
                backtrack: Optional[module])
4.2 Error Handlers
4.3 Backtracking
5 Evaluation
In formulating and evaluating LM Assertions, we consider the following hypotheses:
  1. H1
    LM Assertions facilitate automating self-correction and refinement for arbitrary LM pipelines by showing past outputs and error messages to the LM.
  2. H2
    Self-refinement guided by LM Assertions can also enable LM pipelines to improve downstream application performance.
  3. H3
    When used with DSPy teleprompters to compile few-shot examples, LM Assertions help bootstrap more complex examples to deliver better performance for the compiled LM pipelines.
  4. H4
    Enabling LM Assertions for compilation can collect and optimize demonstrations explicitly for error-correcting behavior (for the Retry modules), which will further improve the ability of the LM application to self-refine.
In this version of our report, we focus on H1 and H2. We leave the investigation for the remaining hypothesis for future versions of our paper. In this report, we explore the hypotheses using two related question-answering (QA) tasks and their corresponding DSPy programs. First, we explore the popular multi-hop QA task and evaluate the effectiveness of LM Assertions at self-refining search queries for better retrieval recall as well as overall answer correctness. Next, we present results on an expanded version of the standard multi-hop QA task in a long-form QA setting. Here, the program must produce a long-form answer with citations to the retrieved context information. We evaluate on this task as an application of LM Assertions to help generate a cohesive, well-structured paragraph that is faithful to the retrieved context.
5.1 Metrics
For evaluating each task, we consider two distinct categories of metrics:
  • Extrinsic Quality assesses the impact of LM Assertions on the pipeline’s performance in downstream, task-specific contexts, often ones we cannot assert directly without access to the ground-truth labels. Here, constraints act as guiding principles that indirectly influence the system’s effectiveness in achieving its objectives. By serving as proxies for more complex goals, the constraints provide insights into how their integration can lead to enhancements in downstream application performance, such as improved accuracy in question-answering tasks or the generation of more coherent and contextually accurate long-form answers.
These two metrics will respectively enable us to investigate the hypotheses that LM Assertions can facilitate self-correction and refinement in LM pipelines and that such guided self-refinement can enhance the performance of downstream applications.
5.2 Dataset & Models
For both tasks, we utilize the HotPotQA (Yang et al., 2018) dataset in the open-domain “fullwiki” setting to evaluate our downstream tasks. Since the HotPotQA test set is not publicly accessible, we reserve the official validation set for testing. We then partition the official training set into subsets: 70% for training and 30% for validation. We only focus on examples labeled as “hard” within the dataset to align with the criteria marked by the official validation and test sets. For training and development sets, we sample 300 examples each. Herein, we report development set scores to avoid repeated evaluation on the test set. We will include expanded test results in a future version of this report.
For retrieval, we use the official Wikipedia 2017 “abstracts” dump of HotPotQA using a ColBERTv2 (Santhanam et al., 2021) retriever. We test the program using OpenAI’s gpt-3.5-turbo (Brown et al., 2020) language model with max_tokens=500 and temperature=0.7 for our experimental setup. For each task, we run two sets of experiments to test the effectiveness of LM Assertions (particularly in the form of Suggest), including both uncompiled and compiled modules. Uncompiled modules ($\mathtt{ZeroShot}$) are zero-shot DSPy modules that infer the response from the questions and context directly. Compiled modules ($\mathtt{FewShot}$) contain additional few-shot demonstrations in the prompt crafted by the DSPy compiler (Khattab et al., 2023).
Table 1: Evaluation of MultiHop Question Answering with HotPotQA. This table presents comparisons across various strategies, including $\mathtt{ZeroShot}$ (uncompiled) and $\mathtt{FewShot}$ (includes few-shot example demonstrations, compiled with DSPy’s BootstrapFewShotWithRandomSearch teleprompter), both with and without assertions ($\mathtt{NoSuggest}$ and $\mathtt{Suggest}$). Metrics measured are Suggestions Passed, Retrieval Score and Answer Correctness (Section 5.3.1)
Configuration Suggestions Passed Retrieval Score Answer Correctness
$\mathtt{ZeroShot}$$\mathtt{NoSuggest}$ 64.3% 34.7% 45.7%
$\mathtt{ZeroShot}$$\mathtt{Suggest}$ 87.3% 39.3% 46.3%
$\mathtt{FewShot}$$\mathtt{NoSuggest}$ 65.0% 40.3% 49.3%
$\mathtt{FewShot}$$\mathtt{Suggest}$ 82.7% 42.0% 50.0%
5.3 Case Study: Multi-Hop QA
5.3.1 Task & Metrics
In this task, we explore complex question answering through the multi-hop program from Khattab et al. (2023). This task involves generating queries related to the given question, iteratively retrieving relevant context from a search index, and synthesizing this information to generate an answer.
Figure 1 shows an implementation of this task in DSPy. As shown, in each hop, a dspy.ChainOfThought module generates a search query for the retriever based on the current accumulated context and the initial question. A dspy.Retrieve then fetches $k$ relevant passages which are aggregated to the context. Once the loop reaches the maximum number of hops, a dspy.ChainOfThought module produces the final answer. With this task and LM pipeline, we aim to produce a robust and thorough reasoning process leading to accurate answers to questions.
We assess intrinsic performance using a Suggestions Passed metric, which quantifies the proportion of suggestions passed across at inference time. For programs with LM Assertions, we measure the last run of the backtracking process (if any). For extrinsic quality, we use (1) Retrieval Score, which quantifies the proportion of correctly retrieved $\mathtt{gold}$ passages, and (2) Answer Correctness, which checks if the generated answer precisely matches the $\mathtt{gold}$ standard answer. Both of these are not possible to use directly at inference time, since they require access to ground-truth labels.
5.3.2 Constraints Specified
We define two simple Suggest statements. First, we suggest that the query generated be short and less than 100 characters, with the goal of reducing imprecise retrieval. Second, we require that a generated query is distinct from previous queries. This discourages the pipeline from retrieving redundant context in each hop. Note that the condition checks for both constraints are implemented using simple Python code. Overall, these constraints enable the self-refinement of the search queries generated by the program.
5.3.3 Evaluation
For this task, we evaluate the MultiHopQA program under various configurations to understand the impact of LM Assertions and few-shot learning techniques. Table 1 shows the results of this evaluation.
Our simplest baseline is the MultiHopQA in a zero-shot, uncompiled configuration ($\mathtt{ZeroShot}$$\mathtt{NoSuggest}$). When we include the mentioned assertions within this zero-shot configuration ($\mathtt{ZeroShot}$$\mathtt{Suggest}$), we see a notable improvement of $35.7\%$ more suggestions passed, meaning the retrieval query generation LM is more likely to generate concise, precise, and distinct queries. Then, we observe $\approx$13.3in the retrieval score and an indirect $\approx$1.3increase in answer correctness due to refined query generation.
We take advantage of optimizations in the DSPy compiling framework, utilizing the BootstrapFewShotWithRandomSearch teleprompter, which automatically generates and integrates few-shot examples doing a random search over a set of candidates within the MultiHopQA program. We include a maximum of 2 bootstrapped few-shot examples to fit within the context window and compile 6 candidates on the devset examples. We apply this evaluation approach to both the vanilla MultiHopQA and the MultiHopQA with Assertions programs. In the non-assertion configuration ($\mathtt{FewShot}$$\mathtt{NoSuggest}$), we see a strong performance gain in retrieval score and correctness, which highlights the efficacy of utilizing few-shot techniques to refine the querying and reasoning processes. We find that combining few-shot learning with assertions ($\mathtt{FewShot}$$\mathtt{Suggest}$) further demonstrates improvements of both $\approx$27.2more passed constraints and $\approx$1.4–4.2on both extrinsic metrics.
Table 2: Evaluation of LongFormQA with Citations Task. This table covers a comparative analysis across different strategies, including $\mathtt{ZeroShot}$ (uncompiled) and $\mathtt{FewShot}$ (includes few-shot example demonstrations, compiled with DSPy’s BootstrapFewShotWithRandomSearch teleprompter), both with and without assertions ($\mathtt{NoSuggest}$ and $\mathtt{Suggest}$). Metrics measured are Citation Faithfulness, Citation Recall, Citation Precision, and Answer Correctness (Section 5.4.1)
Configuration Citation Faithfulness Citation Recall Citation Precision Answer Inclusion
$\mathtt{ZeroShot}$$\mathtt{NoSuggest}$ 76.0% 51.8% 59.4% 66.7%
$\mathtt{ZeroShot}$$\mathtt{Suggest}$ 88.7% 56.3% 64.8% 67.3%
$\mathtt{FewShot}$$\mathtt{NoSuggest}$ 86.3% 62.8% 75.3% 69.3%
$\mathtt{FewShot}$$\mathtt{Suggest}$ 96.0% 62.5% 73.8% 69.3%
5.4 Case Study: LongForm QA
5.4.1 Task & Metrics
In this task, we build on the multi-hop QA (Section 5.3) task by expecting long-form answers to questions that include citations to referenced context.
@@ 5.4.2 Constraints Specified
We introduce more complex suggestions here, unlike in Section 5.3. As a simple initial check, we include a Suggest statement that requires every 1–2 of sentences generated has citations in an intended format. This is checked by a simple Python function citations_check. As a more sophisticated check, we Suggest that the text preceding any citation must be faithful to the cited context, ensuring that the reference text accurately represents the content of the cited information. Since this is a fuzzy condition, we employ a small DSPy program (one that uses the LM) to perform this check. Notably, the robust API design of Suggest allows the user to specify arbitrary expressions as conditional checks, such as an LM call. The goal of this Suggest statement is to ensure that all sentences are appropriately attributed to correct supporting sources.
5.4.3 Evaluation
We extend our analysis to the LongFormQA task to understand how different configurations impact the quality of generated paragraphs with citations to answer questions in terms of citation faithfulness, recall, precision, and answer correctness. We employ a similar evaluation methodology used in the MultiHopQA evaluation (5.3.3).
We start with the simplest baseline: LongFormQA in a zero-shot, uncompiled configuration ($\mathtt{ZeroShot}$$\mathtt{NoSuggest}$). Introducing assertions in this zero-shot setting ($\mathtt{ZeroShot}$$\mathtt{Suggest}$) yields a highly significant $\approx$16.7improvement on the intrinsic citation faithfulness metric, indicating the effect of imposed constraints in maintaining proper citation formatting and referencing. Furthermore, we also observe gains across all extrinsic metrics–recall, precision, and correctness–reflecting the value of effective citation inclusion in enhancing the overall quality of the generated paragraphs in providing composite answers.
Again, we explore enhancements by compiling with few-shot examples in DSPy using a random search (Section 5.3.3). In the non-assertion configuration ($\mathtt{FewShot}$$\mathtt{NoSuggest}$), we observe a performance improvement across metrics as the few-shot demonstrations further refine information retrieval and citations. When including assertions ($\mathtt{FewShot}$$\mathtt{Suggest}$), we continue to see notable gains of $\approx$11.2in citation faithfulness. However, we do see a decline in recall and precision, reflecting some complex interplay between adherence to LM Assertions and optimizing few-shot examples for answer correctness. Yet we observe that introducing computational constraints does not decrease performance on the fundamental downstream metric of answer correctness.
6 Related Work
6.1 Programming with Constraints
Programming with constraints is standard in most programming languages. Popular programming languages like Java (Bartetzko et al., 2001) or Python (Python Software Foundation, 2023) all support expressing assertions as first-class statements to perform runtime checks of certain properties.
However, most runtime checks can only be used to warn the programmer or abort the execution early. Kang et al. (2020) proposed a concept called model assertion, which can be used to monitor the behavior of machine learning models and to improve the quality of a model in training through data collection. However, among the key differences in our work is that LM Assertions abstract automatic self-refinement on violating the assertions. Our Assert and Suggest constructs can backtrack an LM pipeline to retry a failing module. On retry, additional information and feedback are automatically injected into the prompt, guiding the LM program to make better predictions to pass assertions eventually. As described in Sec 5, LM Assertions can additionally be used to inform improved compilation of DSPy programs, though we leave fully specifying and testing this to future versions of our report.
6.2 Self-Refinement and Correction
By integrating Python-style assertions, we ensure programmers can clearly express computational constraints on DSPy programs and assert desired program behavior. These declarative constraints can then be leveraged by the DSPy compiler and runtime in extensible and powerful ways to abstract and generalize notions of self-refinement and DSPy’s capabilities for prompt optimization. We report on initial evaluation of an implementation that does so in this work. Such self-refinement of LLMs (Madaan et al., 2023; Shridhar et al., 2023) is central to this approach in making DSPy autonomous and context-aware (Tyen et al., 2023). Enforcing methodologies of iterative refinement using error feedback (Xu et al., 2023) and utilizing reasoning capabilities through presenting past generations and feedback for correction (Qiu et al., 2023) resonates with the objective of DSPy assertions.
6.3 Program Synthesis and Sketching
Classical program synthesis generates programs based on a given specification, involving a search for candidate programs that match the specification’s structure (Pnueli & Rosner, 1989; Gulwani et al., 2017). In this space, particularly relevant to our work are approaches such as Sketching (Solar-Lezama, 2008) and Template-based synthesis (Srivastava et al., 2013). In sketching (Solar-Lezama, 2009), the programmer specifies a high-level structure (“sketch”) with placeholders (“holes”) for the synthesis engine to complete. The sketch includes assertions, and the correctness requirement is that these assertions hold for all inputs within the bound specified by the synthesizer. We note that the specification assertions in Sketching correspond to the proposed LM Assertions (Assert and Suggest). Similarly, a synthesizer parallels an LM pipeline framework, like the DSPy compiler, striving to compile an optimized prompt.
6.4 Factual Consistency and Citations
This work evaluates LM Assertions on long-form paragraph generation with citations to answer questions (Gao et al., 2023). This task is inspired by several studies with the objective of ensuring factual consistency and citation quality from LLM outputs. Existing frameworks (Min et al., 2023) highlight the complexity of evaluating factual faithfulness in natural language generation. Proposed methodologies like TrueTeacher (Gekhman et al., 2023) and Chain-of-Verification (Dhuliawala et al., 2023) demonstrate LLM-based factual consistency evaluations, which align with the assertion-based methodology in DSPy for ensuring text-to-citation faithfulness. Furthermore, benchmarks for automatic attribution evaluation (Yue et al., 2023) and verifying citations for text generations reflect direct applications for the motivation of this paper, focused on showing the efficacy of asserting the presence and respective accuracy of citing references.
7 Conclusion
We have introduced LM Assertions, a new construct for expressing arbitrary computational constraints on the behavior of LMs when used as building blocks of larger programs. We integrate LM Assertions into the DSPy (Khattab et al., 2023) programming model, define runtime retry semantics, and an implementation for them that abstracts and generalizes LM self-refinement approaches to arbitrary steps in arbitrary pipelines. We also discuss several other mechanisms that our LM Assertion constructs can use to inform DSPy compilation into higher-quality prompts that reduce the assertion failure rates. Our evaluations show substantial gains on two case studies, reporting both intrinsic (i.e., assertion-specific) and extrinsic (i.e., downstream) task metrics. By enabling DSPy programs to autonomously backtrack and self-correct, we hope we have opened avenues for building more reliable LM programs at higher levels of abstraction than was previously possible.
Acknowledgments
We thank Heather Miller, Peter Zhong, and Haoze He for helpful discussions and feedback. This work was partially supported by IBM as a founding member of the Stanford Institute for Human-Centered Artificial Intelligence (HAI), Oracle, Virtusa, and Cigna Healthcare. It was also partially supported by an HAI Azure compute grant. This research was supported in part by affiliate members and other supporters of the Stanford DAWN project–Facebook, Google, and VMware—as well as the NSF under CAREER grant CNS-1651570. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.

### 02 Feb 2024 v2

@@ Appendix B Case Studies
B.1 LongFormQA
B.1.1 Task
In this task, we build on the Multi-Hop QA (Figure 1) task by expecting long-form answers to questions that include citations to referenced context.
@@ B.1.3 Constraints Specified
As a simple initial check, we include a Suggest statement that requires every 1–2 of sentences generated has citations in an intended format. This is checked by a simple Python function citations_check. As a more sophisticated check, we Suggest that the text preceding any citation must be faithful to the cited context, ensuring that the reference text accurately represents the content of the cited information. Since this is a fuzzy condition, we employ a small DSPy program (one that uses the LM) to perform this check. Notably, the robust API design of Suggest allows the user to specify arbitrary expressions as conditional checks, such as an LM call. The goal of this Suggest statement is to ensure that all sentences are appropriately attributed to correct supporting sources.
B.2 QuizGen
B.2.1 Task
We introduce a new task stemming from the HotPotQA dataset in turning questions from the dataset into quiz questions by generating possible answer choices for the question in a JSON format.
class QuizChoiceGenerationWithAssertions(dspy.Module):
    def __init__(self):
        super().__init__()
        self.generate_choices = dspy.ChainOfThought("question, correct_answer, number_of_choices -> answer_choices") #has specified instruction to guide inputs -> outputs
    def forward(self, question, answer):
        choice_string = self.generate_choices(question=question, correct_answer=answer, number_of_choices=number_of_choices).answer_choices
        dspy.Suggest(format_checker(choice_string), "The format of the answer choices should be in JSON format. Please revise accordingly.")
        dspy.Suggest(is_correct_answer_included(answer, choice_string), "The answer choices do not include the correct answer to the question. Please revise accordingly.")
        plausibility_question = "Are the distractors in the answer choices plausible and not easily identifiable as incorrect?"
        plausibility_assessment = dspy.Predict("question, answer_choices, assessment_question -> assessment_answer")(question=question, answer_choices=choice_string, assessment_question=plausibility_question)
        dspy.Suggest(is_plausibility_yes(plausibility_assessment.assessment_answer), "The answer choices are not plausible distractors or are too easily identifiable as incorrect. Please revise to provide more challenging and plausible distractors.")
        return dspy.Prediction(choices = choice_string)
Figure 4: DSPy program with LM Assertions for quiz question choice generation. We introduce 3 suggestions: (1) asserting JSON format; (2) correct answer is included; (3) plausible distractor choices are present.
This task is represented by a very simple program in DSPy with a dspy.ChainOfThought module that generates a set of answer choices based on a defined question-answer pair and a specified number of choices. To ensure well-defined quiz questions, we aim for this task to adhere to consistent formatting and offer a set of plausible distractor answer choices alongside the actual correct answer to the question.
B.2.2 Metrics
We assess the task’s intrinsic performance across the following metrics: (1) Valid Formatting; (2) Correct Answer Inclusion; and (3) Choices’ Plausibility.
We verify consistent formatting by parsing the generated answer choices and checking their consistency to maintain JSON formatting of key-value pairs.
We similarly ensure that the outputted answer choices include the correct answer corresponding to the respective question from the HotPotQA dataset.
For determining the plausibility of the distractor choices, we build a DSPy program that relies on the LM to assess the quality of the answer choice questions. This relies on the inputs: question, generated answer choices, and the assessment question we provide: Are the distractors in the answer choices plausible and not easily identifiable as incorrect? This plausibility verification then outputs an assessment answer of whether the distractors are plausible or not.
For the extrinsic metric, we define a composite scoring metric that considers the intrinsic metrics above. The metric imposes that the conditions of valid formatting and correct answer inclusion are met, thereby ensuring valid quiz questions. When this case is met for the generated answer choices, we return an average score over all three of the intrinsic metrics. If either of these conditions is not met, the score defaults to 0.
B.2.3 Constraints Specified
For the simple check of Valid Formatting, we include a Suggest statement that requires the format of the answer choices to be in JSON format. This is checked by a simple Python function format_checker.
Similarly, we verify Correct Answer Inclusion with the Suggest statement that indicates if the answer choices do not include the correct answer. This is checked by a simple Python function is_correct_answer_included.
To verify the plausibility of the answer choices to reflect strong distractor choices alongside the correct choice, we employ the Suggest statement to indicate if the answer choices are not plausible distractors or are too easily identifiable as incorrect. With a DSPy program in place to assess the choices, this Suggest statement ensures that all of the answer choices are plausible distractors.
B.3 TweetGen
B.3.1 Task
We introduce another new task derived from the HotPotQA dataset in generating tweets to answer questions.
class TweetGenerationWithAssertions(dspy.Module):
    def __init__(self):
        super().__init__()
        self.generate_tweet = dspy.ChainOfThought("question, context -> tweet") #has specified instruction to guide inputs -> outputs
    def forward(self, question, answer):
        context = []
        generate_query = [dspy.ChainOfThought("context, question -> query") for _ in range(2)]
        retrieve = dspy.Retrieve(k=3)
        for hop in range(2):
            query = generate_query[hop](context=context, question=question).query
            passages = retrieve(query).passages
            context = deduplicate(context + passages)
        generated_tweet = self.generate_tweet(question=question, context=context).tweet
        dspy.Suggest(has_no_hashtags(generated_tweet), f"Please revise the tweet to remove hashtag phrases following it.")
        dspy.Suggest(is_within_length_limit(generated_tweet, 280), f"Please ensure the tweet is within {280} characters.")
        dspy.Suggest(has_correct_answer(generated_tweet, answer), "The tweet does not include the correct answer to the question. Please revise accordingly.")
        engaging_question = "Does the assessed text make for a self-contained, engaging tweet? Say no if it is not engaging."
        engaging_assessment = dspy.Predict("context, assessed_text, assessment_question -> assessment_answer")(context=context, assessed_text=generated_tweet, assessment_question=engaging_question)
        dspy.Suggest(is_assessment_yes(engaging_assessment.assessment_answer), "The text is not engaging enough. Please revise to make it more captivating.")
        faithful_question = "Is the assessed text grounded in the context? Say no if it includes significant facts not in the context."
        faithful_assessment = dspy.Predict("context, assessed_text, assessment_question -> assessment_answer")(context=’N/A’, assessed_text=generated_tweet, assessment_question=faithful_question)
        dspy.Suggest(is_assessment_yes(faithful_assessment.assessment_answer), "The text contains unfaithful elements or significant facts not in the context. Please revise for accuracy.")
        return dspy.Prediction(generated_tweet=generated_tweet, context=context)
Figure 5: DSPy program with LM Assertions for tweet generation. We introduce 5 suggestions: (1) asserting no hashtags; (2) correct answer is included; (3) tweet is within character limit; (4) tweet is engaging; (5) tweet is faithful to context.
This task mirrors the MultiHopQA task with the addition of a dspy.ChainOfThought module layer to utilize the retrieved context and corresponding question to generate a tweet that effectively answers the question. We aim for the task to ensure the tweet not only answers the question but is engaging to the reader and faithful to its relevant context.
B.3.2 Metrics
We assess the task’s intrinsic performance across various metrics: (1) No Hashtags; (2) Correct Answer Inclusion; (3) Within Length; (4) Engaging; (5) Faithful.
We impose an intrinsic constraint to ensure none of the tweets have hashtags, ensuring all tweets maintain a consistent tweeting style.
As we do with QuizChoiceGeneration, we ensure the outputted tweet includes the correct answer corresponding to the respective question from the HotPotQA dataset.
We also ensure that the generated tweet adheres to a character count limit of 280 characters to model sample tweet behavior.
For determining the engagement of the tweet, we build a DSPy program that relies on the LM to assess this. This relies on the inputs: question, context, generated tweet, and the assessment question we provide: Does the assessed text make for a self-contained, engaging tweet? This verification outputs its assessment of whether the tweet is engaging in relation to its corresponding question and retrieved context.
We perform a similar assessment for the tweet’s faithfulness, with the simple modification to the assessment question: Is the assessed text grounded in the context?
For the extrinsic metric, we define a composite scoring metric that considers all of the intrinsic metrics above. The metric imposes that the most relevant intrinsic conditions of a well-formed tweet are met, particularly if the tweet contains the correct answer to the question and is within the tweeting character limit. When this case is met for the generated answer choices, we return an average score over all five of the intrinsic metrics. If either of these conditions is not met, the score defaults to 0.
B.3.3 Constraints Specified
To verify that the tweet contains no hashtags, we include a Suggest statement that requires the tweet to be generated without any hashtag phrases. This is checked by a simple Python function through regex checks in has_no_hashtags.
To verify the generated tweet adheres to the character limits, we impose this through the Suggest statement to ensure that the tweet is under the specified character limit, which we specify as 280 in our experiments. This is checked by a simple Python function is_within_length_limit.
Similarly, we verify Correct Answer Inclusion with the Suggest statement that indicates if the answer choices do not include the correct answer. This is checked by a simple Python function has_correct_answer.
To verify the engagement level of the generated tweet, we employ the Suggest statement to simply indicate whether the tweet is engaging enough as determined by the LM and DSPy program in place to assess engagement.
We conduct a similar approach for faithfulness as well, checking for the tweet’s faithfulness to its retrieved context.
Appendix C Impact on Using Different LLM Instructions
We explore comparative tests in the specified instructions for the case studies mentioned above. We differentiate between a primitive instruction that aims to simply specify a task’s objective and a complete instruction that accounts for the respective intrinsic and extrinsic metric measured for the task. These tests are conducted specifically on the TweetGen and QuizGen tasks which encompass more complex metrics. Our experiments on the complete instructions are presented in Figure 2 while we demonstrate our results on the primitive instructions below.
C.1 TweetGen
Primitive instruction: "Generate a tweet that effectively answers a question."
Complete instruction with metrics accounted for: "Generate an engaging tweet that effectively answers a question staying faithful to the context, is less than 280 characters, and has no hashtags."
C.2 QuizGen
Primitive instruction: "Generate answer choices for the specified question."
Complete instruction with metrics accounted for: "Generate answer choices in JSON format that include the correct answer and plausible distractors for the specified question."
C.3 Discussion
Based on these experiments on primitive instructions, we discovered that when the baseline pipeline only has access to high-level and generic instructions, it is almost impossible for the pipeline to follow the underlying constraints. For example, in QuizGen, the baseline $\mathtt{Vanilla}$ strategy only generates 2.8% of quizzes with Correct JSON format and 2.6% of quizzes that contains the correct answer.
However, for our assertion-driven optimization enabled pipelines, the performance on primitive experiments are still comparable to the counter-part with complex instructions. This indicates that model pipelines with LM Assertions and assertion-driven optimizations are less sensitive to instructions, requiring less effort on manual prompt tuning.
TweetGen w/ Primitive Instructions (Dev/Test)
Strategy No "#" Has Answer Concise Engaging Faithful Quality
$\mathtt{Vanilla}$ 3.3 / 3.0 53.7 / 48.2 96.3 / 97.0 35.7 / 36.4 80.0 / 81.2 33.7 / 30.4
$\mathtt{Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$ 49.3 / 49.6 50.3 / 41.8 92.0 / 92.4 45.3 / 41.0 72.3 / 74.0 34.3 / 27.8
$\mathtt{Compile}$ 0.0 / 0.2 55.7 / 46.2 100 / 99.6 47.3 / 46.6 78.3 / 76.8 36.7 / 30.8
$\mathtt{Compile}$ $\mathtt{w/}$ $\mathtt{Assert}$ 98.7 / 97.4 55.0 / 45.8 99.3 / 99.0 1.3 / 2.6 65.3 / 70.0 40.4 / 34.3
$\mathtt{C\texttt{+}Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$ 41.3 / 41.0 55.7 / 48.2 94.7 / 93.8 54.3 / 60.2 76.7 / 81.2 40.3 / 35.0
QuizGen w/ Primitive Instructions (Dev/Test)
Strategy Correct JSON Has Answer Citation Precision Validity
$\mathtt{Vanilla}$ 1.3 / 2.8 1.3 / 2.6 61.3 / 61.8 1.2 / 2.3
$\mathtt{Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$ 91.7 / 93.4 73.3 / 72.6 75.0 / 69.8 69.8 / 68.0
$\mathtt{Compile}$ 100 / 100 94.3 / 89.8 72.7 / 67.4 85.4 / 80.1
$\mathtt{Compile}$ $\mathtt{w/}$ $\mathtt{Assert}$ 100 / 100 95.7 / 91.4 63.0 / 57.0 83.7 / 78.5
$\mathtt{C\texttt{+}Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$ 100 / 100 93.3 / 89.4 73.7 / 67.8 85.8 / 81.1
Figure 6: Evaluation of TweetGen and QuizGen task using the primitive instruction. The LM pipeline configuration are described in Table 1. For each task, we use the same LM pipeline program except for the LM Assertions. Extrinsic metrics (downstream application performance) are highlighted in grey. For each metric, higher is always better. The highest value in each column is bold.
Appendix D Interesting Examples of Correlated Constraints
This section showcases selected examples that demonstrate potential conflicts between Assertions specified in the TweetGen task. Each set of examples highlights instances of two configurations that produce outputs to achieve success on the TweetGen Metrics (Section B.3.2) and face the nuanced challenge of balancing these constraints in execution.
Example 1
Both results accurately include the correct answer. However, $\mathtt{Compile}$ $\mathtt{w/}$ $\mathtt{Assert}$ does not pass the Engaging constraint despite the specified assertion but ensures the No "#" constraint is met whereas $\mathtt{C\texttt{+}Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$ violates the No "#" constraint while satisfying the Engaging constraint.
Question: What was the name of the treaty that made Hungary a landlocked state which contained the Kolozsvar Ghetto?
$\mathtt{Compile}$ $\mathtt{w/}$ $\mathtt{Assert}$: The Treaty of Trianon made Hungary landlocked and led to the existence of the Kolozsvar Ghetto. This lesser-known ghetto was located in Kolozsvár, Kingdom of Hungary (now Cluj-Napoca, Romania).
$\mathtt{C\texttt{+}Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$: The Treaty of Trianon, signed in 1920, reshaped Hungaryś destiny, leading to its landlocked status and the emergence of the Kolozsvar Ghetto. Uncover the gripping tale of this pivotal moment in history and its lasting effects. #HistoryUnveiled
Answer: Treaty of Trianon
Example 2
Both results do not include the correct answer. However, both configurations produce engaging outputs. Yet $\mathtt{C\texttt{+}Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$ does not conform to the No "#" constraint while $\mathtt{Compile}$ $\mathtt{w/}$ $\mathtt{Assert}$ does.
Question: Which American car rental company is also a member of the Association of Car Rental Industry Sytems Standards? $\mathtt{Compile}$ $\mathtt{w/}$ $\mathtt{Assert}$: "Enterprise, a leading American car rental company, is also a proud member of the Association of Car Rental Industry Systems Standards. Rent with confidence and enjoy a seamless experience with Enterprise!"
$\mathtt{C\texttt{+}Infer}$ $\mathtt{w/}$ $\mathtt{Assert}$: "Looking for a car rental company that meets industry standards? Look no further than ACRISS member Enterprise! With their commitment to excellence, you can trust them for a seamless rental experience. Get ready for a smooth ride! #CarRental #Enterprise"
Answer: Budget Rent a Car